1b9mA T0108  gshxqaeilltlklqqklfadprrisllkhialsgsisqgakdagisyksawdainexnqlsehilvera ...................................................................... 10 1b9mA T0108  tggkggggavltrygqrliqlydllaqiqqkafdvlsdddalplnsllaaisrfs--------------- .......................................................MGASGQYVRARIKGA 20        30        40        50        60        70        80 1b9mA T0108  ---------LQTSARNQWFGTITARDHDDVQQHVDVLLADGKTRLKVAITAQSGARLG-----LDEGKEV YYATPVDPVTNQPTAPKDFSSGFWDFNDGTTQGFGVNPDSPITAINVENANNALKISNLNSKGSNDLSEG 90          100       110       120       130       140       150 1b9mA T0108  LILLKapwVGITQDEAVAQNADNQLPGIISHIERGAEQCEVLXALPDGQ-------TLCATVPVNEATSL NFWAN...VRISADIWGQSINIYGDTKLTMDVIAPTPVNVSIAAIPQSSTHGWGNPTRARIVWTNNFVAQ 160       170       180       190       200 1b9mA T0108  QQGQNVTAYF--------------------------------------------nadsviiatlc TDGTYKATLTISTNDSPNFNTIATDAADSVVTNMILFVGSNSDNISLDNIKFTK...........

Figure 1: Comparison of candidate 1b9mA-T0108-local.pw.a2m (lines) to reference T0108-1b9mA-2track-local.pw.a2m (letters). Reference alignment has 126 matching columns, Candidate alignment has 0 matching columns. 0 were used (for template shift+target shift) to compute shift score. The shift score is 9999.000000, with a min shift of 0, max shift of 999999, and mean of 0.000000. Optimal subalignment has 0 matching columns, with 258 paired matches. 0 were used (for template shift+target shift) to compute shift score. Optimum shift score is 0.000000, with a min shift of 0, max shift of 999999, and mean of 9999.000000. Optimum achieved by removing the 258 columns indicated with dotted lines from the candidate alignment.